A small body rolls horizontally of the edge of the table top which is 1.20 m high. It strikes at the floor at point 1.52 m horizontally from the table edge.

a.) How long is the ball in the air?
b.) What is its speed at the instant it leaves the table?
1
Expert's answer
2019-08-14T09:44:18-0400

(a) Time of motion of the ball

"t=\sqrt{\frac{2H}{g}}=\sqrt{\frac{2\times 1.20\:\rm{m}}{9.81\:\rm{m/s^2}}}=0.49\:\rm{s}"

(b) The initial speed of the ball

"v=\frac{L}{t}=\frac{1.52\:\rm{m}}{0.49\:\rm{s}}=3.10\:\rm{m/s}"